singly ionized helium atom is hydrogen like in the sense that a solitary electron revolves around a positively charged nucleus. If the energy of this electron in its first orbit(n=1) is -54.4ev what will be it energy in the first excited state?
 
plzzz reply.........URGENT

Grade:11

1 Answers

Sunil Kumar FP
9 years ago
energy is given by=-13.6*z^2/n^2= -13.6 *4/4=-13.6ev
here n=2 z=2=atomic number
thanks and regards
sunil kr
askITian faculty